A rotating merry-go-round makes one complete revolution in 5.0s. Assume the wheel is moving with a constant angular velocity. What is the linear speed of a child seated 1.1m from the center? What is her acceleration (give components)?
Answer #1


T= 5 s
R=1.1 m

V= Rw= 1.1 * 2*pi/T
v=1.3816

acceleration a = V^2/R=1.7352 m/s^2


as it is moving with constant speed it has only radial component of acceleratio..tangential componenet of acceleration is zero.